OpenStudy (anonymous):

graph the equation. identify the conic sequence, any lines of symmetry, and the domain and range.

OpenStudy (anonymous):

\[x^2+y^2=4\]

OpenStudy (anonymous):

help @whpalmer4

OpenStudy (whpalmer4):

that's a conic section, not sequence

OpenStudy (whpalmer4):

do you know the general formula for a circle?

OpenStudy (whpalmer4):

or how to describe the points that make up a circle?

OpenStudy (anonymous):

um not really and lol my bad

OpenStudy (whpalmer4):

Well, a circle is the set of points that are all the same distance from its center, correct?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay. Glad we agree :-) 

OpenStudy (whpalmer4):

 Pythagorean theorem says that \[x^2+y^2 = r^2\]right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think

OpenStudy (whpalmer4):

Okay, so that's where the equation for a circle comes from, assuming the center of the circle is at the origin (0,0). If it isn't, then the form becomes \[(x-h)^2 +(y-k)^2 = r^2\]where the circle's center is at \((h,k)\). I think you can easily convince yourself that that more general equation simplifies to the one you have when \((h,k)=(0,0)\).

OpenStudy (anonymous):

yeah